Calculus II, Spring 2009 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 05

Assigned Mar 26, 2009.

1. Section 10.6: Problems 15, 17, 31, 35, 37.

2. Section 10.8: Problems 22, 31, 32, 24, 28, 35, 37.

3. Section 11.7: problems 3, 7, 9, 27, 29, 31, 35, 43.